### v1

# OLMES: A Standard for Language Model Evaluations

Abstract
Progress in AI is often demonstrated by new models claiming improved performance on tasks measuring model capabilities. Evaluating language models in particular is challenging, as small changes to how a model is evaluated on a task can lead to large changes in measured performance. There is no common standard setup, so different models are evaluated on the same tasks in different ways, leading to claims about which models perform best not being reproducible. We propose OLMES, a completely documented, practical, open standard for reproducible LLM evaluations. In developing this standard, we identify and review the varying factors in evaluation practices adopted by the community – such as details of prompt formatting, choice of in-context examples, probability normalizations, and task formulation. In particular, OLMES supports meaningful comparisons between smaller base models that require the unnatural “cloze” formulation of multiple-choice questions against larger models that can utilize the original formulation. OLMES includes well-considered recommendations guided by results from existing literature as well as new experiments investigating open questions. [^1] All prompts, examples, and code used for OLMES are available at https://allenai.org/data/olmes
1 Introduction
Scientific credibility in AI rests on fair, reproducible comparisons between models. Many current AI models, such as pretrained large language models (LLMs), are generalist models capable of performing downstream tasks they were not specifically trained on (OpenAI,, 2024; Bommasani et al.,, 2022). When evaluating LLMs on such tasks, there are many choices in how the task is presented to the model and how the model outputs are interpreted before scoring (Gao,, 2021; Biderman et al.,, 2024; Liang et al.,, 2023). There is currently no standard way to decide on these choices, and they can have a huge impact on model performance, with some recent papers claiming as much as an 80% difference in accuracy on a given task just from varying formatting and in-context examples (Sclar et al.,, 2023). These choices are rarely reported with enough information to actually reproduce, so when a team of ML practitioners releases a new model it is often impossible for them to compare against previously-reported results.
Efforts like the Holistic Evaluation of Language Models (HELM) benchmark (Liang et al.,, 2023) and the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023) strive towards standardizing LM evaluations. While the same setup is used to evaluate many models ensuring consistency and reproducibility, the rationale behind prompt formatting, use of in-content examples, normalization techniques, and task formulation are not always clearly documented and thus not consistently followed by other researchers in subsequent work (Touvron et al., 2023b, ; Biderman et al.,, 2023; Jiang et al.,, 2023; Groeneveld et al.,, 2024; AI@Meta,, 2024).
We highlight two problems in the field today. (1) Releasing a new model and comparing against previously reported results is flawed unless the previous work explicitly described their full evaluation setup, and then that setup is followed in the new work. Unfortunately, such results are not comparable to previous work that used a different format. Importantly, the key goal of comparing models on the same downstream task is to measure their relative strengths. This is a real problem, as we show in Table 1, which illustrates how several models’ published performance on the ARC-Challenge (Clark et al.,, 2018) and OpenbookQA (Mihaylov et al.,, 2018) tasks varies in the literature. For instance, looking at Ref1 for ARC-Challenge, we would conclude that Llama2-13B and Llama3-8B are performing similarly, but Ref6 reveals there is likely a gap of over 10% between them. For OpenbookQA, on Falcon-7B, there is a discrepancy of more than 25% between Ref2 and Ref8, which only makes sense if one considers the very different task formulations. (2) Some papers are likely reporting over-estimates of performance by optimizing the prompt format for their model, while others are not. In fact, we see evidence that this might be a community-wide problem in a recent effort from HELM (Mai and Liang,, 2024). When evaluating a suite of models using a single format and comparing against each model’s self-reported result, they find significant differences in performance. In particular, their findings suggest that models with higher self-reported performance were more likely to have lower scores in the reproduction, suggesting some over-estimating of model performance (see Figure 4 in Appendix).
Table 1: Scores reported in different references for LLM performances on ARC-Challenge and OpenbookQA. Scores indicated with are using multiple-choice formulation (MCF) rather than “cloze” formulation (CF) (see Section 2.1 for definitions). Entries with “?” denote either undocumented or mixed approaches across models. Different references use different evaluation setups, some of which are not fully specified, so conclusions about which models perform best are not reproducible.
ARC-Challenge Evaluations: OpenbookQA Evaluations:
[…]
MPT-7B 47.7 42.6 46.5 45.7 51.4 48.6 52.4
RPJ-INCITE-7B 46.3 42.8 45.3 49.4 49.0
Falcon-7B 47.9 42.4 44.5 47.5 49.7 51.6 44.6 53.0 $26.0^{\dagger}$ 55.2
Mistral-7B 60.0 55.5 54.9 $78.6^{\dagger}$ 52.2 $77.6^{\dagger}$ $80.6^{\dagger}$
Llama2-7B 53.1 45.9 43.2 45.9 48.5 $53.7^{\dagger}$ 54.2 58.6 58.6 48.4 58.6 $54.4^{\dagger}$ 57.8
Llama2-13B 59.4 49.4 48.8 49.4 $67.6^{\dagger}$ $67.3^{\dagger}$ 57.0 57.0 57.0 $63.4^{\dagger}$ $65.4^{\dagger}$
Llama3-8B 60.2 $78.6^{\dagger}$ $79.3^{\dagger}$ $76.6^{\dagger}$ $77.2^{\dagger}$
Num shots 25 0 0 0 0 25 5 0 0 0 0 5 5
Curated shots No No Yes No Yes
Formulation RC RC RC? RC RC MC MC/RC RC RC RC RC MC MC/RC
Normalization char char ? char? pmi none none/pmi pmi pmi? pmi pmi? none none/pmi
Ref Reference citation Ref Reference citation
Ref1 HF Open LLM Leaderboard (Beeching et al.,, 2023) Ref5 OLMo paper (Groeneveld et al.,, 2024)
Ref2 Llama2 paper (Touvron et al., 2023a, ) Ref6 Llama3 model card (AI@Meta,, 2024)
Ref3 Mistral 7B (Jiang et al.,, 2023) Ref7 Gemma paper (Gemma Team et al.,, 2024)
Ref4 Falcon paper (Almazrouei et al.,, 2023) Ref8 HELM Lite Leaderboard (Liang et al.,, 2023)
To address these problems, we present OLMES (Open Language Model Evaluation Standard), a standard to improve the transparency and reproducibility of language model evaluation from a practical point of view, removing ambiguity in how a final performance metric is obtained when evaluating a model on a dataset. OLMES can be applied to evaluation during the model development process, leaderboards, papers, providing justified recommendations on data sampling, how to format instances, the choice of in-context examples, probability normalization, and task formulation.
Importantly, OLMES is:
  • Reproducible: OLMES specifies all details of the evaluations, from processing datasets to presenting the task to model, to processing models’ outputs, so there are no ambiguities in the evaluation procedure.
  • Practical: OLMES makes practical decisions in use of computation resources for easy adoption by the community.
  • Documented: Each decision in the standard is documented with justifications by applying principles from existing studies and performing experiments to investigate open questions.
  • Open: We release all prompts and code, along with the rationales behind the choices made in OLMES, for subsequent work to follow and build upon by extending the same principles to any new task and model.
Since OLMES is a documented, practical, open evaluation standard, it is straightforward to adopt in publicly available, well-maintained evaluation code bases like the Eleuther LM Evaluation Harness (Gao et al.,, 2023; Biderman et al.,, 2024) and HELM (Liang et al.,, 2023). When used by model developers and other researchers, OLMES will help unify evaluation practices in the field. All prompts, examples, and code used for OLMES can be found at https://allenai.org/data/olmes.
2 Experimental setup
2.1 Multiple-choice QA and LLM evaluation
Multiple-choice question answering (MCQA) tasks present a compelling way of evaluating models and humans alike, due to the ease of scoring (whether the correct answer is chosen out of the given options) and the allowed flexibility in the domain and complexity of the questions. Such tasks form a large, essential part of LM evaluations and are be the focus in OLMES. There are generally two ways to formulate these tasks.
MCF (Multiple-choice formulation):
presenting answer choices indicated by labels and scoring prediction of answer labels, just like how MCQA is posed to humans. Here is an example of MCQA from ARC-Easy (Clark et al.,, 2018), a dataset of real grade-school level science questions:
Question: Earth’s core is primarily composed of which of the following materials? (A) basalt (B) iron (C) magma (D) quartz
Answer: (B)
CF (Completion/cloze formulation):
scoring each answer choice separately using LM token probabilities. The MCF format is not compatible with the pure language modeling nature of generating the next token. Therefore, the CF format was introduced by when evaluating the GPT-3 model (Brown et al.,, 2020). They found that it was possible to elicit much better performance using a “cloze” completion version of the task, where the model is shown a prompt like:
Question: Earth’s core is primarily composed of which of the following materials?
Answer: <answer>,
where each answer choice is separately substituted in for <answer>.
Then the LM probability of the answer choice tokens are used to rank the choices and predict an answer. This formulation has natural limitations, such as not being able to properly address cases where one answer choice is “none of the above”.
2.2 Targeted tasks
Table 2: OLMES details on tasks, with our standardized choices of dataset split, number of instances to use (along with total number if sampling was used), and which CF normalization scheme to use (see Section 3.3). Column #C shows the number of answer choices (ARC-Challenge and ARC-Easy have a few instances with 3 or 5 answer choices).
task split #C # inst (total) CF norm reference
ARC-Challenge (ARC_C) Test $4^{\dagger}$ 1172 pmi (Clark et al.,, 2018)
ARC-Easy (ARC_C) Test $4^{\dagger}$ 1000 (2376) char (Clark et al.,, 2018)
BoolQ Val 2 1000 (3270) none (Clark et al.,, 2019)
CommonsenseQA (CSQA) Val 5 1221 pmi (Talmor et al.,, 2019)
HellaSwag (HSwag) Val 4 1000 (10042) char (Zellers et al.,, 2019)
MMLU Test 4 14042 char (Hendrycks et al.,, 2021)
OpenbookQA (OBQA) Test 4 500 pmi (Mihaylov et al.,, 2018)
PIQA Val 2 1000 (1838) char (Bisk et al.,, 2020)
Social IQa (SIQA) Val 3 1000 (1954) char (Sap et al.,, 2019)
WinoGrande (WinoG) Val 2 1267 none (Sakaguchi et al.,, 2020)
We select and implement standards for 10 popular benchmark MCQA tasks, see Table 2 for the list. The list covers tasks that are frequently used in the community’s evaluation practices, such as the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023), Llama papers (Touvron et al., 2023a, ; Touvron et al., 2023b, ; AI@Meta,, 2024), HELM (Liang et al.,, 2023), and the OLMo evaluation suite (Groeneveld et al.,, 2024). This selection includes questions on science, various types of commonsense, factual knowledge, and covers a range of topics (MMLU alone covers 57 subjects), of varying difficulty.
2.3 Selection of models
We develop OLMES based on a selection of 15 diverse, openly available pretrained LLMs, covering a range of sizes from 1B to 70B – Pythia-1B, Pythia-6.7B (Biderman et al.,, 2023), OLMo-1B, OLMo-7B, OLMo-1.7-7B (Groeneveld et al.,, 2024), TinyLlama-1.1B (Zhang et al.,, 2024), StableLM2-1.6B (Bellagente et al.,, 2024), RPJ-INCITE-7B (Together Computer,, 2023), MPT-7b (MosaicML,, 2023), Falcon-7B (Almazrouei et al.,, 2023), Llama2-7B, Llama2-13B (Touvron et al., 2023b, ), Mistral-7B-v0.1 (Jiang et al.,, 2023), Llama3-8B, Llama3-70B (AI@Meta,, 2024). This reflects our goal of providing an evaluation standard that suits a range of model capabilities, with the flexibility to apply the same methodology during model development as well as when comparing final powerful models.
3 Standardizing variations in evaluation
To evaluate a model on a dataset, there are a variety of decisions that have to be made to get a final accuracy metric of that model on that dataset. These include:
  • How to format dataset instances? (Section 3.1)
  • Which few-shot examples to use? (Section 3.2)
  • How to normalize LLM probabilities for CF? (Section 3.3)
  • What task formulation to use, MCF or CF? (Section 3.4)
  • Other implementation choices impacting results (Appendix A)
Below we enumerate key variations in these steps, and justify the choices made in OLMES (listed in Table 2) to standardize these steps, leaving some of the details for the Appendix.
3.1 How to format dataset instances?
Each MCQA dataset includes a set of fields used to specify an instance, such as question, answer choices, and perhaps a context for the question. When formatting an instance as a prompt to an LLM, many different choices have been made in the literature. This includes simple choices like "Question:" vs "Q:" as question prefix (varying even within a paper, e.g., Brown et al., (2020)), or formatting the answer labels (e.g., "A." (Touvron et al., 2023a, ), "(A)" (Nori et al.,, 2023), "<mc>A</mc>" (Anthropic,, 2024), etc). There is also a choice of whether or not to provide a general instruction (e.g., common for MMLU (Hendrycks et al.,, 2021), sometimes done for OpenbookQA (Almazrouei et al.,, 2023)).
Instance formatting. OLMES uses a consistent "Question: <question>" prefix and "Answer:" suffix in formatting the datasets. This clarifies the question-answering task in a natural way, without relying on verbose instruction understanding. The three exceptions are listed and explained here. For PIQA, we use "Goal: <goal>" as the prefix instead to be consistent with the original semantics of the dataset. In the case of MCF, for HellaSwag, we skip the question prefix and instead add "Choose the best continuation:" before presenting the continuation options, and for WinoGrande we use the prefix "Fill in the blank:" to align with the task. For HellaSwag and WinoGrande, where the CF answer string is simply a language continuation, we remove such prefixes and suffixes for the CF evaluation so that the task is closer to pure language modeling.
[…]
Sampling. Following existing LM evaluation standardization efforts (Liang et al.,, 2023; Beeching et al.,, 2023), OLMES uses the test split of a dataset if the labels are publicly available, otherwise the validation split. If the dataset has more than 1500 instances, we sample 1000 instances to evaluate on, [^2] Sampling uses a specific random seed in Python: Random(1234).sample(all_instances, 1000) similar to HELM (Liang et al.,, 2023) which caps evaluation instances at 1000.333https://crfm-helm.readthedocs.io/en/latest/reproducing_leaderboards/ Note that the potential extra statistical signal from more instances would generally be dominated by other sources of score variations, like prompt formatting, so this is a practical consideration to avoid unnecessary computation resources. See Table 2 for details on splits and sampling used in OLMES.
3.2 Which few-shot examples to use?
Popularized by Brown et al., (2020), it is customary to provide examples of the task to the model through few-shot examples, as this is an effective and universal way to convey a task to an LLM. For example, the MMLU task (Hendrycks et al.,, 2021) originally came with a fixed 5-shot prompt which is generally used in evaluation (Beeching et al.,, 2023; Gemma Team et al.,, 2024; Jiang et al.,, 2023; Touvron et al., 2023b, ; AI@Meta,, 2024) resulting in a more reproducible results than many other tasks. [^4] Sometimes sampled examples are used also for MMLU (MosaicML,, 2024). Even for MMLU, noticeable discrepancies have been found, due to other differences in prompt formatting (Mai and Liang,, 2024)). For other tasks, both the number of shots and the way in which they are sampled have varied in different evaluation setups. For example, to evaluate on HellaSwag, Beeching et al., (2023) sampled 10-shot whereas HELM Liang et al., (2023) uses 0-shot; within Beeching et al., (2023), a range of 25-shot, 10-shot, 5-shot was sampled for ARC-Challenge, HellaSwag and WinoGrande respectively.
OLMES standardizes a manually curated 5-shots prompt for each task (from its training set), ensuring that the examples are of good quality and cover the label space in a balanced way (e.g., avoiding 4 A’s and 1 B among the 5 answers). [^5] More details on curating the examples can be found in Appendix F. Restricting to 5 in-context examples helps limit computational overhead, similar to HELM (Liang et al.,, 2023). Analysis suggests that going beyond 5 shots generally does not provide meaningful differences in scores (Brown et al.,, 2020; Barton,, 2024). These manually curated shots for each task can be downloaded from https://allenai.org/data/olmes.
3.3 How to normalize LLM probabilities for CF?
When using the cloze/completion formulation (CF) for multiple-choice questions, the LLM returns $P(a_{i}|q)$ to denote the probability for an answer choice $a_{i}$ given a question prompt $q$. Ranking solely based on the probability may heavily favor shorter answers with fewer tokens. To work around this issue, different normalization methods have been used in the literature, which we categorize below:
  • none: $\ln(P(a_{i}|q))$
  • token: $\ln(P(a_{i}|q))/\operatorname{num\_tokens}(a_{i})$, which normalizes the log-probability by the number of tokens in the answer (Brown et al.,, 2020)
  • character: $\ln(P(a_{i}|q))/\operatorname{num\_characters}(a_{i})$, which normalizes the log-probability by the number of characters in the answer, used by Llama models (Touvron et al., 2023a, ) and Eleuther AI LM Harness (Gao et al.,, 2023; Biderman et al.,, 2024).
  • pmi: $\ln(P(a_{i}|q)/P(a_{i}|u))$ where $u=$"Answer:" is the unconditional prompt, which normalizes by dividing by LLM probability of the same answer string without the presence of the question (like "Answer: <answer_string>"). This can be considered a form of pointwise-mutual-information (PMI) and was explored further in other works (Holtzman et al.,, 2021).
Efforts like Liang et al., (2023); Gao et al., (2023); Biderman et al., (2024) compare and support comparisons of different normalization approaches, leaving it an open question as to how to make a decision. See Appendix C.2 for further discussions around different normalizations.
To choose a normalization scheme in OLMES, we evaluate the models on each dataset using CF, comparing the effect of the 4 normalization techniques. Table 3 shows for each task, how often each normalization is empirically the best across the 15 models. Detailed scores per model are in Appendix C.2.
OLMES specifies the “pmi” normalization for ARC-Challenge, CommonsenseQA, and OpenbookQA. The answer choices in these datasets tend to contain unexpected words or phrases that are less likely for models to generate (e.g., “Whirlpool bath” compared to “Bathtub”). The pmi normalization adjusts for this by taking into account the a priori likelihood of the answers. This is consistent with other findings (Holtzman et al.,, 2021) and some existing evaluation practices, e.g., Brown et al., (2020) selectively uses this normalization for ARC and OpenbookQA, and Touvron et al., 2023a ; Touvron et al., 2023b for OpenbookQA. Computing the extra unconditional likelihood incurs some computation overhead, thus OLMES avoids this normalization for other datasets where there is no strong empirical or theoretical reason to choose this approach.
OLMES specifies the “character” normalization for ARC-Easy, HellaSwag, PIQA, Social IQa and MMLU. Based on our experiments, it is empirically the normalization technique that gives the best scores [^6] Tie for PIQA, and second-best for MMLU. for these datasets, and less computationally expensive than the “pmi” normalization. It also has the advantage (unlike the “token” normalization) of already being implemented (as acc_norm) in the Eleuther LM Evaluation Harness, where it is generally available for multiple-choice tasks (Gao et al.,, 2023; Gao,, 2021). It is also used in the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023) for ARC-Challenge and HellaSwag, in Touvron et al., 2023a ; Touvron et al., 2023b ’s evaluations as the default, (with select datasets as exceptions), as well as reported in various works like Biderman et al., (2023); Almazrouei et al., (2023).
OLMES specifies the “none” normalization for BoolQ and WinoGrande. In BoolQ the only answer choices are “yes” or “no” which are single tokens, therefore no length normalization is needed. Note that for some models, the “character” normalization has slightly better performance on BoolQ (see Table 9), an accidental side effect of “yes” having one more character than “no”. One could argue that the pmi normalization is appropriate as it counters any existing bias in the model for “yes” vs “no”, but we argue that models should be capable of producing such common words (also indicated in the 5-shot examples) without any such corrections. Finally, WinoGrande is a special case in that the continuations are identical (and the prompts vary), so the choice of normalization does not matter and we simply use the “none” normalization.
In general, we observe little difference between the OLMES recommendation and the empirically best (“oracle”) normalization for each task and model, see “diff best” column in Table 3 (Table 8 in Appendix C.2 has more details).
3.4 What task formulation to use, MCF or CF?
As LLMs have gotten stronger, the MCQA task formats have gradually changed from CF to MCF. For instance, ARC-Challenge is often evaluated using the CF approach (Touvron et al., 2023a, ; Touvron et al., 2023b, ; Almazrouei et al.,, 2023; Beeching et al.,, 2023), but has switched to MCF for stronger models like OpenAI, (2024); AI@Meta, (2024), appearing with identical names like “25-shot ARC-Challenge”. As an example, AI@Meta, (2024) reports a 25-shot MCF ARC-Challenge score for the Llama-3 8B model of 78.6% vs 59.2% for the 25-shot CF on the Hugging Face Open LLM Leaderboard. As performance on a multiple-choice task gets closer to 100%, the CF approach lags behind and give significantly less signal about a model’s actual performance. On the other hand, MMLU is almost exclusively evaluated using the MCF approach, which often results in near-random performance for weaker models (Beeching et al.,, 2023).
In OLMES, we argue that the CF formulation provides a useful evaluation of task knowledge for models that have not yet acquired the skill of answering multiple-choice questions using MCF. On the other hand, MCF is a more realistic formulation for models that can “understand” this format, yielding higher and more representative scores (Robinson et al.,, 2023; OpenAI,, 2024). See Appendix C.1 for further discussion.
Figure 1: Performance on MMLU validation set during the training of OLMo-1.7-7B model. During early training, there is good signal from CF while MCF is random. Around 400B tokens, the model starts gaining the ability on the MCF format, becoming a stronger signal than CF.
Table 3: Summary of CF normalization comparisons. “win percentage” shows how often each normalization was best across the 15 models. “diff best” shows that there is in general minimal difference between the OLMES normalization and the empirically optimal normalization for each task (difference out of 100%).
win percentage diff
task none char tok pmi best OLMES
ARC_C 0.0 33.3 0.0 66.7 0.2 pmi
ARC_E 6.7 86.7 6.7 0.0 0.1 char
BoolQ 46.7 46.7 0.0 6.7 1.1 none
CSQA 6.7 33.3 6.7 53.3 0.6 pmi
HSwag 0.0 100.0 0.0 0.0 0.0 char
MMLU 0.0 46.7 0.0 53.3 0.4 char
OBQA 0.0 0.0 0.0 100.0 0.0 pmi
PIQA 6.7 46.7 46.7 0.0 0.2 char
SIQA 0.0 86.7 6.7 6.7 0.1 char
WinoG 100.0 0.0 0.0 0.0 0.0 none
We can see an explicit example of a model acquiring the “understanding” of MCF during training in Figure 1, showing the OLMo-1.7-7B model (Groeneveld et al.,, 2024; AI2 blog,, 2024) evaluated on the MMLU validation set in both CF and MCF variations. The plot suggests that model starts learning the MCF task format after about 400 billion training tokens, so in early training CF provides a better signal, while MCF is significantly better in late training.
To further study this phenomenon, we evaluate the CF and MCF versions on each task and model. [^7] More detailed numbers can be found in Tables 5 and 6 in the Appendix B Figure 2 shows for each task, the CF and MCF performance for the 15 models ordered along the x-axis by overall performance on all tasks. For instance, on ARC-Challenge, we see a clear distinction where the weakest 8 models have near-random performance on the MCF version of the task, yet above random when using CF which offers a better signal to the relative strength of models. For the stronger models, the MCF version clearly outscores the CF version, and is a much better representation of task performance compared to the flatter trends using CF (for Llama3-70B the MCF score is 93.7% (6.3% error) while the CF score is just 69.0% (31% error), a nearly 5x difference in error rate!).
Figure 2: Comparing model performance on each task for MCF vs CF. The 15 models are ordered along the x-axis by overall performance across all 10 tasks. In general, CF is needed to elicit a non-random signal from the weaker models, while stronger models can take advantage of MCF for a more accurate assessment.
A similar pattern can be seen across other tasks in Figure 2, where the stronger models show performance using MCF either exceeding CF (like ARC-Easy, OpenbookQA, MMLU, Social IQa, CommonsenseQA, and PIQA) or at least catching up to it (HellaSwag, WinoGrande, BoolQ). [^8] Appendix C.2.1 provides further discussion.
In OLMES, we standardize to evaluate each model using both the MCF and CF formulations, and the best performing one is used. This allows for meaningful comparison of task evaluation numbers over a range of models, from the smaller, weaker base models which can only deal with the CF (where MCF scores hovering around random baseline), to the stronger models which can report more accurate performance using the MCF (where CF provides less clear signal).
4 OLMES: Summary and results
OLMES includes the following elements, justified in detail above:
  • Use test set when available, otherwise validation. Sample 1000 instances if more than 1500
  • Use specified, exact prompt format (Section 3.1)
  • Use fixed, curated 5-shot examples (Section 3.2)
  • Use prescribed probability normalization for CF (Section  3.3)
  • Evaluate with both MCF and CF, use the best result (Section 3.4)
  • Follow recommendations for all other evaluation details (in Appendix A)
Table 4 reports scores for different models on popular benchmark tasks, following our evaluation standard, OLMES.
Table 4: Reproducible performance scores across models and tasks using OLMES, providing robust, meaningful comparisons across a wide range of models and tasks. indicates use of the MCF score.
model ARC_C ARC_E BoolQ CSQA HSwag MMLU OBQA PIQA SIQA WinoG average
Pythia-1B 31.4 63.4 $56.8^{\dagger}$ 50.9 48.0 31.1 40.4 68.9 46.4 52.7 49.0
OLMo-1B 38.6 68.3 51.3 62.2 65.2 33.4 47.6 74.1 51.5 59.3 55.1
TinyLlama-1.1B 38.1 69.5 63.6 61.1 60.8 33.6 45.0 71.7 50.4 60.1 55.4
Pythia-6.7B 44.6 72.6 68.7 62.1 66.1 37.7 50.4 74.9 51.7 62.3 59.1
RPJ-INCITE-7B 45.3 78.8 72.0 69.2 72.8 40.1 49.0 75.9 56.6 68.0 62.8
StableLM2-1.6B $50.6^{\dagger}$ 75.3 82.3 $70.4^{\dagger}$ 70.3 $40.4^{\dagger}$ $56.6^{\dagger}$ 75.6 $64.3^{\dagger}$ 65.7 65.1
OLMo-7B 46.4 78.9 78.7 70.8 78.1 40.5 55.8 78.5 56.5 68.5 65.3
MPT-7b 45.7 78.0 82.4 70.9 79.6 40.6 52.4 79.2 57.4 70.2 65.6
Falcon-7B 49.7 80.6 78.2 73.4 79.0 42.1 55.2 79.0 60.1 71.3 66.9
Llama2-7B 54.2 84.0 86.1 74.2 78.9 $46.2^{\dagger}$ 57.8 77.5 59.6 71.7 69.0
Llama2-13B $67.3^{\dagger}$ 85.9 86.7 74.0 83.9 $55.8^{\dagger}$ $65.4^{\dagger}$ 80.2 $65.9^{\dagger}$ 74.9 74.0
OLMo-1.7-7B $66.9^{\dagger}$ $83.6^{\dagger}$ 85.9 $85.8^{\dagger}$ 80.1 $54.4^{\dagger}$ $68.6^{\dagger}$ 80.3 $76.1^{\dagger}$ 73.6 75.5
Llama3-8B $79.3^{\dagger}$ $92.4^{\dagger}$ 87.5 $73.9^{\dagger}$ 81.8 $66.6^{\dagger}$ $77.2^{\dagger}$ 81.6 $70.2^{\dagger}$ 76.2 78.7
Mistral-7B-v0.1 $78.6^{\dagger}$ $90.8^{\dagger}$ 89.3 $72.4^{\dagger}$ 83.0 $64.0^{\dagger}$ $80.6^{\dagger}$ 82.8 $71.3^{\dagger}$ 77.9 79.1
Llama3-70B $93.7^{\dagger}$ $97.7^{\dagger}$ $91.7^{\dagger}$ $83.2^{\dagger}$ 89.5 $79.8^{\dagger}$ $93.4^{\dagger}$ $91.6^{\dagger}$ $78.9^{\dagger}$ 84.1 88.4
5 Related work
With model releases, performance on popular datasets is used to gauge the progress achieved e.g., OpenAI, (2024)t showing superhuman performance on benchmarks like MMLU. Such evaluation also guides the community efforts towards understanding and sharing findings on what it takes to build a strong model (Touvron et al., 2023a ; Touvron et al., 2023b ; Biderman et al., (2023); Almazrouei et al., (2023); MosaicML, (2023); Jiang et al., (2023); Gemma Team et al., (2024); Groeneveld et al., (2024) inter alia).
However, given a model and a dataset, even for the frequently used datasets, there are varied practices in how accuracy on them is measured. Various work has shown that model evaluations are vulnerable to differences such as option position changes in multiple-choice questions (Zheng et al.,, 2024; Li et al.,, 2024), choice symbols, re-ordering of answer options, changing number of answer options (Wang et al.,, 2024), and task formulation (Alzahrani et al.,, 2024; Robinson et al.,, 2023; Khatun and Brown,, 2024; Wiegreffe et al.,, 2023). Even minor formatting changes can cause large, generally arbitrary, score variations (Sclar et al.,, 2023). The Holistic Evaluation of Language Models (HELM) benchmark (Liang et al.,, 2023), the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023), Mosaic Eval Gauntlet (Barton,, 2024) and Eleuther AI LM Harness (Gao et al.,, 2023; Biderman et al.,, 2024) present efforts toward greater transparency and reproducibility of LM evaluations. Liang et al., (2023); Gao et al., (2023); Biderman et al., (2024) describe and provide support for variability factors, presenting them as open choices to researchers and users. Even when specifications are made, the rationale is not always documented and thus not followed by others in subsequent work (see Table 1).
6 Discussion
By identifying and reviewing common evaluation practices in the community, and performing experiments to investigate open questions, we present OLMES – an open, documented, reproducible, and practical evaluation standard.
Future work and limitations. Our efforts for future work include adding more tasks to OLMES, covering tasks beyond MCQA such as generative tasks and chain-of-thought prompting. This encompasses standardizing how answers are extracted for evaluation, and for chat models how to split the prompt into messages. We also encourage the community to contribute to OLMES, extending the principles of OLMES to new tasks.
We hope OLMES will be a meaningful step towards standardizing evaluations, and incorporated into evaluation code bases for broad usage. This will facilitate robust and simplified comparisons of model performances, both for researchers during model training and development and for developers choosing models to build upon.
Acknowledgments and Disclosure of Funding
In creating this evaluation standard, OLMES, we build on top of the various previous efforts on language model evaluation in the community – including previous work on language model evaluation standardization, the many open research reports disclosing how evaluation on LMs have been done, and the datasets that made OLMES possible, which we explicitly cite and acknowledge in our paper.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] In Discussion section.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [N/A]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] URL at the end of Introduction section.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix E.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] In Section 2.
    2. (b)
      Did you mention the license of the assets? [N/A] All models and datasets used can be research for research purposes.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] In Appendix F.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A OLMES: implementation details
There are other important details that go into a fully specified evaluation result, and we enumerate the choices made in OLMES here:
  • For MMLU: use macro average (over 57 tasks) rather than micro average (over 14042 instances), following AI@Meta, (2024). This better represents the diversity of fields in the dataset, although in practice it does not generally make a big difference (see Figure 7).
  • When a model requires it, make sure to add the appropriate <bos> token at start of prompt (e.g., Gemma (Gemma Team et al.,, 2024)).
  • When using the “character” normalization for CF, include the leading space in the calculation of answer length.
  • Restrict all inputs (with completions) to 2048 tokens for consistency across models [^9] For current tasks this is only exhausted for a few MMLU instances.
  • Use the default model precision when evaluating (i.e., avoid options like load_in_8bit unless it produces identical results)
  • OLMES uses the standard approach of two newlines to separate each in-context example.
  • Other than the original instruction line for MMLU Hendrycks et al., (2021), we do not add any extra instructions. This is in view of previous work finding the subject information from instructions makes little changes to model ranking (Alzahrani et al.,, 2024), and to reduce additional sources of variation in the prompt.
Note that computational details, like batch size and type/state of GPU, can affect floating point operations such that answer choice decisions can flip if they are very close. This is hard to avoid unless consider “ties” when answers are sufficiently close in confidence.
Appendix B Detailed CF and MCF task scores
Tables 5 and  6 present detailed scores across all tasks, with both MCF and CF results (using the OLMES recommendations for CF normalization).
Appendix C Further details on variations
In this appendix we discuss further details on how LLM evaluations can vary and the choices made in OLMES.
C.1 Task formulation details
LLM evaluations started out using the CF approach for many tasks (Brown et al.,, 2020; Du et al.,, 2022; Smith et al.,, 2022; Chowdhery et al.,, 2022; Lieber et al.,, 2021), which is a more reasonable option for weaker models that struggle with the more natural MCF (Khatun and Brown,, 2024). The task formulation only very recently and gradually switched to the MCF approach when it became clear that the model could utilize it, producing higher scores (Robinson et al.,, 2023; OpenAI,, 2024; AI@Meta,, 2024).
The HELM study (Liang et al.,, 2023) included comparisons between the MCF (“joint”) and CF (“separate”) approaches, finding that certain models can really benefit from the MCF approach, although among the models in the original study it was really only the Anthropic-LM v4-s3 (52B) model which could take full advantage of it.
C.2 CF normalization details
Tables  9,  10 and  11 show detailed comparisons of CF normalization on different models, for the various tasks.
Unlike in MCF, where the evaluation metric involves just scoring the log-likelihood corresponding to the answer choice label (i.e., A/B/C/…), there is a choice of log-likelihood normalization (“none”, “per token”, “per character” or “pmi”) for CF as detailed in Section 3.3.
When evaluating the GPT-3 model(Brown et al.,, 2020), they worked around this issue by normalizing the log-probability by the number of tokens in the answer (similar to how loss is computed during training). They also noted that for a few datasets, it worked markedly better to instead “normalize” by dividing by LLM probability of the same answer string without the presence of the question (usually by just having a generic prefix like "Answer: <answer_string>"). This can be considered a form of pointwise-mutual-information (PMI) and was explored further in other works (Holtzman et al.,, 2021).
The Eleuther LM Evaluation Harness (Gao et al.,, 2023; Biderman et al.,, 2024) and some subsequent evaluations (e.g., the Llama models (Touvron et al., 2023a, )) have also used “per answer character” normalization, using the argumentation (Gao,, 2021; Biderman et al.,, 2024), that normalizing per token is problematic since it depends on the tokenizer. Since the purpose of the normalization is simply to rank the answer choices within themselves (keeping model and tokenizer fixed), this does not seem like a relevant argument, and indeed a normalization which changes the probability of “yes” vs “no” simply because the “no” token has fewer characters seem problematic. In practice, for tasks where answers either relatively long or similar in length, there are minor differences between these two length normalization.
The HELM study (Liang et al.,, 2023) included comparisons between these normalization approaches for a number of tasks and models (using the terms “separate” and “separate calibrated” for “token” and “pmi” respectively), eventually settling on a default choice for each, not unlike the choices in the GPT-3 report (Brown et al.,, 2020). The Eleuther LM Evaluation Harness generally reports two metrics for each multiple-choice task: acc (using the “none” normalization) and acc_norm (using the “character” normalization).
Table 5: Comparing MCF and CF scores on each task (part 1). Weaker models at the top of the table have near-random MCF scores, while for stronger models at the bottom, the MCF score provides a better assessment than the CF score.
ARC_C ARC_E BoolQ CSQA HSwag MMLU
model MCF CF MCF CF MCF CF MCF CF MCF CF MCF CF
Pythia-1B 24.1 31.4 24.0 63.4 56.8 56.6 21.0 50.9 23.6 48.0 26.5 31.1
OLMo-1B 25.3 38.6 25.4 68.3 37.9 51.3 20.2 62.2 24.6 65.2 26.6 33.4
TinyLlama-1.1B 26.4 38.1 24.3 69.5 60.7 63.6 17.9 61.1 26.2 60.8 26.2 33.6
Pythia-6.7B 26.6 44.6 24.9 72.6 64.0 68.7 20.5 62.1 24.3 66.1 25.4 37.7
RPJ-INCITE-7B 28.1 45.3 25.1 78.8 64.2 72.0 19.7 69.2 23.5 72.8 29.0 40.1
StableLM2-1.6B 50.6 47.3 69.6 75.3 60.1 82.3 70.4 68.2 52.4 70.3 40.4 37.1
OLMo-7B 27.2 46.4 27.0 78.9 67.5 78.7 20.8 70.8 25.0 78.1 28.3 40.5
MPT-7b 27.9 45.7 27.5 78.0 44.6 82.4 20.9 70.9 26.4 79.6 30.0 40.6
Falcon-7B 27.2 49.7 25.3 80.6 48.1 78.2 20.2 73.4 27.7 79.0 28.0 42.1
Llama2-7B 52.6 54.2 70.6 84.0 57.5 86.1 59.2 74.2 41.4 78.9 46.2 44.4
Llama2-13B 67.3 56.2 85.0 85.9 77.8 86.7 68.1 74.0 62.4 83.9 55.8 47.6
OLMo-1.7-7B 66.9 51.2 83.6 81.5 82.0 85.9 85.8 70.4 50.0 80.1 54.4 42.4
Llama3-8B 79.3 57.1 92.4 86.6 84.8 87.5 73.9 69.9 63.8 81.8 66.6 51.1
Mistral-7B-v0.1 78.6 59.6 90.8 86.8 87.2 89.3 72.4 72.3 71.5 83.0 64.0 50.3
Llama3-70B 93.7 69.0 97.7 89.6 91.7 91.2 83.2 75.8 89.1 89.5 79.8 60.7
Table 6: Comparing MCF and CF scores on each task (part 2), along with overall averages. The “max” average corresponds to the OLMES score, taking the best of MCF and CF for each task.
OBQA PIQA SIQA WinoG average scores
model MCF CF MCF CF MCF CF MCF CF MCF CF all max
Pythia-1B 26.0 40.4 52.2 68.9 33.5 46.4 50.4 52.7 33.8 49.0 41.4 49.0
OLMo-1B 28.0 47.6 50.6 74.1 32.8 51.5 51.1 59.3 32.3 55.1 43.7 55.1
TinyLlama-1.1B 25.6 45.0 50.2 71.7 34.9 50.4 50.0 60.1 34.2 55.4 44.8 55.4
Pythia-6.7B 26.2 50.4 51.2 74.9 33.5 51.7 49.6 62.3 34.6 59.1 46.9 59.1
RPJ-INCITE-7B 22.6 49.0 53.5 75.9 33.7 56.6 52.1 68.0 35.1 62.8 49.0 62.8
StableLM2-1.6B 56.6 51.0 62.8 75.6 64.3 61.1 53.5 65.7 58.1 63.4 60.7 65.1
OLMo-7B 27.0 55.8 57.2 78.5 35.1 56.5 50.4 68.5 36.6 65.3 50.9 65.3
MPT-7b 29.6 52.4 53.8 79.2 34.4 57.4 51.1 70.2 34.6 65.6 50.1 65.6
Falcon-7B 27.8 55.2 50.5 79.0 33.9 60.1 49.0 71.3 33.8 66.9 50.3 66.9
Llama2-7B 54.8 57.8 63.2 77.5 58.7 59.6 52.4 71.7 55.7 68.8 62.2 69.0
Llama2-13B 65.4 60.8 74.0 80.2 65.9 63.6 56.1 74.9 67.8 71.4 69.6 74.0
OLMo-1.7-7B 68.6 59.8 65.6 80.3 76.1 54.9 56.2 73.6 68.9 68.0 68.5 75.5
Llama3-8B 77.2 56.2 77.3 81.6 70.2 62.6 61.6 76.2 74.7 71.0 72.9 78.7
Mistral-7B-v0.1 80.6 61.0 79.0 82.8 71.3 63.0 59.8 77.9 75.5 72.6 74.1 79.1
Llama3-70B 93.4 69.0 91.6 83.1 78.9 65.6 79.6 84.1 87.9 77.8 82.8 88.4
Table 7: Macro vs micro average scores on MMLU, where macro average is over the 57 tasks and micro average is over the 14042 individual questions. In general there are small differences between the two.
model MCF-macro MCF-micro CF-macro CF-micro
Pythia-6.7B 25.4 25.2 37.7 37.5
TinyLlama-1.1B 26.2 25.7 33.6 33.5
Pythia-1B 26.5 26.4 31.1 31.2
OLMo-1B 26.6 26.3 33.4 33.6
Falcon-7B 28.0 27.7 42.1 41.9
OLMo-7B 28.3 28.3 40.5 40.7
RPJ-INCITE-7B 29.0 28.4 40.1 40.1
MPT-7b 30.0 29.3 40.6 40.6
StableLM2-1.6B 40.4 39.6 37.1 37.0
Llama2-7B 46.2 45.5 44.4 44.3
OLMo-1.7-7B 54.4 52.8 42.4 42.4
Llama2-13B 55.8 55.5 47.6 47.1
Mistral-7B-v0.1 64.0 63.0 50.3 49.8
Llama3-8B 66.6 65.4 51.1 50.8
Llama3-70B 79.8 79.2 60.7 60.5
Table 8: Normalization details, showing that our recommendations are not only supported by reasoning using principles behind the normalization but also close to the empirically best normalization that lets you get the highest accuracy for each model on each task (see “diff” columns).
ARC_C ARC_E BoolQ CSQA HSwag MMLU OBQA PIQA SIQA
model pmi diff char diff none diff pmi diff char diff char diff pmi diff char diff char diff
Pythia-1B 31.4 0.0 63.4 0.0 56.6 4.5 50.9 0.0 48.0 0.0 31.1 1.2 40.4 0.0 68.9 1.4 46.4 0.0
OLMo-1B 38.6 0.0 68.3 0.2 51.3 4.7 62.2 0.0 65.2 0.0 33.4 0.8 47.6 0.0 74.1 0.0 51.5 0.0
TinyLlama-1.1B 38.1 0.0 69.5 0.0 63.6 2.2 61.1 0.0 60.8 0.0 33.6 0.9 45.0 0.0 71.7 0.6 50.4 0.0
Pythia-6.7B 44.6 0.0 72.6 0.0 68.7 0.0 62.1 0.2 66.1 0.0 37.7 0.2 50.4 0.0 74.9 0.0 51.7 1.1
RPJ-INCITE-7B 45.3 0.0 78.8 0.0 72.0 2.5 69.2 0.2 72.8 0.0 40.1 0.8 49.0 0.0 75.9 0.1 56.6 0.0
MPT-7b 45.7 0.6 78.0 0.0 82.4 0.0 70.9 0.0 79.6 0.0 40.6 0.0 52.4 0.0 79.2 0.0 57.4 0.0
Falcon-7B 49.7 0.0 80.6 0.0 78.2 0.6 73.4 0.0 79.0 0.0 42.1 0.0 55.2 0.0 79.0 0.2 60.1 0.0
OLMo-7B 46.4 0.0 78.9 0.0 78.7 0.0 70.8 0.0 78.1 0.0 40.5 0.1 55.8 0.0 78.5 0.8 56.5 0.0
StableLM2-1.6B 47.3 0.0 75.3 0.0 82.3 0.0 68.2 0.0 70.3 0.0 37.1 1.5 51.0 0.0 75.6 0.3 61.1 0.0
Llama2-7B 54.2 0.0 84.0 0.0 86.1 0.0 74.2 0.0 78.9 0.0 44.4 0.4 57.8 0.0 77.5 0.2 59.6 0.0
OLMo-1.7-7B 51.2 0.0 81.5 0.0 85.9 0.0 70.4 1.1 80.1 0.0 42.4 0.0 59.8 0.0 80.3 0.0 54.9 0.8
Llama2-13B 56.2 0.9 85.9 0.0 86.7 1.5 74.0 0.0 83.9 0.0 47.6 0.0 60.8 0.0 80.2 0.0 63.6 0.0
Llama3-8B 57.1 1.3 86.6 0.0 87.5 0.3 69.9 4.3 81.8 0.0 51.1 0.0 56.2 0.0 81.6 0.0 62.6 0.0
Mistral-7B-v0.1 59.6 0.6 86.8 0.0 89.3 0.0 72.3 2.1 83.0 0.0 50.3 0.0 61.0 0.0 82.8 0.0 63.0 0.0
Llama3-70B 69.0 0.0 89.6 0.8 91.2 0.5 75.8 1.3 89.5 0.0 60.7 0.0 69.0 0.0 83.1 0.1 65.6 0.0
Table 9: Comparing CF normalization schemes (part 1).
ARC_C ARC_E BoolQ
model none char tok pmi best none char tok pmi best none char tok pmi best
Pythia-1B 26.1 28.4 29.0 31.4 pmi 61.9 63.4 60.9 56.5 char 56.6 61.1 56.6 41.0 char
OLMo-1B 32.9 34.4 34.7 38.6 pmi 68.5 68.3 65.8 60.2 none 51.3 56.0 51.3 42.3 char
TinyLlama-1.1B 31.5 34.1 32.2 38.1 pmi 68.6 69.5 64.4 60.4 char 63.6 65.8 63.6 53.6 char
Pythia-6.7B 36.3 39.5 39.0 44.6 pmi 71.4 72.6 70.0 64.1 char 68.7 66.9 68.7 47.6 none
RPJ-INCITE-7B 40.3 43.5 42.9 45.3 pmi 76.1 78.8 75.9 70.1 char 72.0 74.5 72.0 72.4 char
MPT-7b 41.7 46.3 44.7 45.7 char 76.3 78.0 76.2 68.5 char 82.4 79.9 82.4 76.7 none
Falcon-7B 41.6 47.4 47.6 49.7 pmi 77.0 80.6 78.3 69.8 char 78.2 78.8 78.2 77.6 char
OLMo-7B 41.6 45.5 45.0 46.4 pmi 76.7 78.9 77.4 69.6 char 78.7 77.7 78.7 78.6 none
StableLM2-1.6B 42.2 44.3 44.9 47.3 pmi 73.3 75.3 74.4 70.0 char 82.3 82.0 82.3 76.1 none
Llama2-7B 48.4 52.0 50.2 54.2 pmi 81.4 84.0 81.0 74.7 char 86.1 85.6 86.1 80.5 none
OLMo-1.7-7B 45.5 49.3 48.5 51.2 pmi 79.2 81.5 79.7 71.1 char 85.9 83.8 85.9 85.6 none
Llama2-13B 52.4 57.1 54.2 56.2 char 83.9 85.9 82.8 77.6 char 86.7 88.2 86.7 77.5 char
Llama3-8B 53.6 58.4 56.8 57.1 char 85.8 86.6 85.8 76.6 char 87.5 87.8 87.5 67.0 char
Mistral-7B-v0.1 56.1 60.2 58.9 59.6 char 84.7 86.8 84.6 78.6 char 89.3 89.1 89.3 89.2 none
Llama3-70B 65.7 69.0 67.7 69.0 char 89.7 89.6 90.4 82.6 tok 91.2 90.4 91.2 91.7 pmi
average scores 43.7 47.3 46.4 49.0 NA 77.0 78.7 76.5 70.0 NA 77.4 77.8 77.4 70.5 NA
win percentage 0.0 33.3 0.0 66.7 pmi 6.7 86.7 6.7 0.0 char 46.7 46.7 0.0 6.7 none
Table 10: Comparing CF normalization schemes (part 2)
CSQA HSwag MMLU
model none char tok pmi best none char tok pmi best none char tok pmi best
Pythia-1B 47.7 50.9 47.3 50.9 char 39.2 48.0 47.8 41.0 char 29.5 31.1 30.8 32.3 pmi
OLMo-1B 56.8 60.0 57.6 62.2 pmi 50.9 65.2 64.1 49.8 char 31.7 33.4 33.3 34.2 pmi
TinyLlama-1.1B 58.9 60.5 55.9 61.1 pmi 46.9 60.8 59.7 48.5 char 31.2 33.6 33.0 34.5 pmi
Pythia-6.7B 59.5 62.2 58.9 62.1 char 50.4 66.1 65.9 53.5 char 34.9 37.7 37.0 37.9 pmi
RPJ-INCITE-7B 67.7 69.4 67.2 69.2 char 55.7 72.8 71.8 60.6 char 37.4 40.1 40.0 40.9 pmi
MPT-7b 69.6 70.3 69.1 70.9 pmi 60.5 79.6 76.5 61.5 char 37.8 40.6 40.1 40.4 char
Falcon-7B 70.0 70.3 69.5 73.4 pmi 60.7 79.0 78.4 60.0 char 39.3 42.1 41.9 42.1 char
OLMo-7B 69.0 70.0 67.9 70.8 pmi 59.3 78.1 76.3 64.2 char 37.9 40.5 40.5 40.6 pmi
StableLM2-1.6B 63.6 66.3 65.6 68.2 pmi 54.7 70.3 69.7 56.4 char 35.2 37.1 37.1 38.6 pmi
Llama2-7B 70.5 72.7 68.4 74.2 pmi 61.9 78.9 77.1 64.4 char 42.0 44.4 43.9 44.8 pmi
OLMo-1.7-7B 71.6 63.5 59.0 70.4 none 61.4 80.1 77.7 65.2 char 39.9 42.4 42.2 41.8 char
Llama2-13B 72.2 72.7 68.4 74.0 pmi 63.7 83.9 81.0 70.3 char 44.3 47.6 46.7 47.1 char
Llama3-8B 72.0 74.2 73.5 69.9 char 62.8 81.8 80.3 71.1 char 47.5 51.1 50.8 49.6 char
Mistral-7B-v0.1 73.1 73.8 74.4 72.3 tok 64.5 83.0 81.0 70.3 char 46.9 50.3 50.0 49.0 char
Llama3-70B 77.1 77.1 77.1 75.8 char 70.3 89.5 87.1 80.8 char 57.2 60.7 60.5 59.4 char
average scores 66.6 67.6 65.3 68.4 NA 57.5 74.5 73.0 61.2 NA 39.5 42.2 41.9 42.2 NA
win percentage 6.7 33.3 6.7 53.3 pmi 0.0 100.0 0.0 0.0 char 0.0 46.7 0.0 53.3 pmi
Table 11: Comparing CF normalization schemes (part 3).
OBQA PIQA SIQA
model none char tok pmi best none char tok pmi best none char tok pmi best
Pythia-1B 20.2 28.6 30.4 40.4 pmi 70.3 68.9 68.8 60.1 none 42.8 46.4 46.0 44.4 char
OLMo-1B 26.0 33.0 38.4 47.6 pmi 73.2 74.1 73.2 59.9 char 45.3 51.5 49.9 47.3 char
TinyLlama-1.1B 24.4 34.8 35.8 45.0 pmi 72.1 71.7 72.3 62.0 tok 45.6 50.4 48.2 48.4 char
Pythia-6.7B 25.8 37.0 37.4 50.4 pmi 74.8 74.9 74.3 63.6 char 48.0 51.7 52.8 49.2 tok
RPJ-INCITE-7B 31.8 40.0 42.8 49.0 pmi 74.9 75.9 76.0 61.9 tok 50.8 56.6 56.0 52.2 char
MPT-7b 31.6 43.8 43.8 52.4 pmi 77.7 79.2 78.1 63.7 char 51.0 57.4 55.9 52.5 char
Falcon-7B 35.2 45.8 44.4 55.2 pmi 78.3 79.0 79.2 63.2 tok 52.9 60.1 57.5 54.4 char
OLMo-7B 33.2 42.8 45.0 55.8 pmi 78.2 78.5 79.3 65.2 tok 50.3 56.5 56.5 52.8 char
StableLM2-1.6B 34.4 41.6 45.2 51.0 pmi 75.2 75.6 75.9 63.6 tok 52.7 61.1 60.7 56.1 char
Llama2-7B 33.8 44.6 45.0 57.8 pmi 76.7 77.5 77.7 62.9 tok 52.6 59.6 58.3 53.6 char
OLMo-1.7-7B 37.2 48.4 49.6 59.8 pmi 78.5 80.3 79.3 66.3 char 53.5 54.9 54.3 55.7 pmi
Llama2-13B 39.2 46.4 48.4 60.8 pmi 78.9 80.2 79.8 66.4 char 56.7 63.6 60.7 56.8 char
Llama3-8B 37.0 47.6 50.0 56.2 pmi 79.7 81.6 81.1 67.5 char 54.6 62.6 60.1 56.4 char
Mistral-7B-v0.1 38.2 48.4 50.0 61.0 pmi 80.8 82.8 81.3 67.4 char 55.6 63.0 60.9 57.5 char
Llama3-70B 47.0 55.0 56.6 69.0 pmi 82.8 83.1 83.2 68.3 tok 59.7 65.6 64.8 57.3 char
average scores 33.0 42.5 44.2 54.1 NA 76.8 77.6 77.3 64.1 NA 51.5 57.4 56.2 53.0 NA
win percentage 0.0 0.0 0.0 100.0 pmi 6.7 46.7 46.7 0.0 char 0.0 86.7 6.7 6.7 char
C.2.1 Tasks that generally prefer CF
HellaSwag and WinoGrande continue to have CF scores higher than MCF scores even for the strongest models that can understand the MCF prompt. This somewhat surprising tendency seems correlated with the fact that these tasks in the CF format are exactly like the language modeling task of finding the most natural continuation of a running piece of text. Judging from the trends in the plot, it would also be interesting to monitor if as even more capable models are developed, the MCF scores will eventually surpass that of the CF scores (given how close they already get to each other).
C.2.2 Hybrid formulation
In CF, overall probability score could be quite misleading since it may heavily favor shorter answers with fewer tokens. Note that this would be different if the answer choices are actually listed before scoring the answer string, then most tokens (after the choice has been disambiguated by the first few tokens) would have probability near one. This “hybrid” formulation has been used in some cases, but usually scores in between the CF and MCF approaches (Wiegreffe et al.,, 2023). However, this hybrid approach is not popular in evaluation standardization efforts like the Open LLM Leaderboard, HELM, or when used to evaluate models during development, so it is not a focus in OLMES.
C.3 Tokenization of MCQA choice labels
[…]
Figure 3: Tokenizer example, showing two examples of tokenizers which need a prefix space before MCQA answer choice labels to represent the choice label and the final answer label using the same token.
> from transformers import AutoTokenizer
> llama_tokenizer = AutoTokenizer.from_pretrained("meta-llama/Llama-2-7b-hf")
> olmo_tokenizer = AutoTokenizer.from_pretrained("allenai/OLMo-1.7-7B-hf")
[…]
> llama_tokenizer.tokenizer(test_string)
[’_What’, ’_is’, ’_’, ’3’, ’+’, ’4’, ’?’, ’<0x0A>’, ’_A’, ’.’, ’_’, ’7’, ’<0x0A>’, ’A’, ’.’, ’_’, ’7’, ’<0x0A>’, ’Answer’, ’:’, ’_A’] > olmo_tokenizer.tokenizer(test_string)
[’What’, ’Ġis’, ’Ġ3’, ’+’, ’4’, ’?’, ’Ċ’, ’ĠA’, ’.’, ’Ġ7’, ’Ċ’, ’A’, ’.’, ’Ġ7’, ’Ċ’, ’Answer’, ’:’, ’ĠA’]
Appendix D HELM Reproduction of MMLU
In Figure 4 we see data taken from HELM’s reproduction of MMLU scores for a variety of models.
Figure 4: Self-reporting overestimates MMLU score compared to reproduction, from https://crfm.stanford.edu/2024/05/01/helm-mmlu.html. Each point corresponds to a model, the x-axis shows self-reported MMLU score, and the y-axis shows the difference between the self-reported score and the reproduced score. Points above the y=0 line have higher self-reported performance than the reproduction; the trend line has a positive slope, indicating that on average, the higher the self-reported score the more they overestimate performance compared to the reproduction.
Appendix E Compute used
The inference on the models evaluated were done on NVIDIA RTX A6000 GPUs. A total of around 400 GPU hours was used.
Appendix F Curation of 5-shot examples: considerations
Procedure for manually curating the few-shot examples:
  • Download the train set from Hugging Face datasets
  • Start from the beginning of the training set, looking at a batch of 10 (i.e., start with first 10)
  • Skip ambiguous instances
  • Skip instances that hint at discrimination or otherwise deemed inappropriate
  • Skip instances if the same label has appeared frequently (e.g., 4 consecutive instances with gold label ‘C’, keep better ones out of those)
  • If instances are grouped/labeled by topic, choose instances to be diverse (e.g. first 3 are all about a certain topic, pick from later ones to ensure diversity)
  • If you end up with less than 7 instances that cover the label space or range of different topics, look at the next batch of 10.
  • Finally, reorder instances to obtain a somewhat balanced output of answer labels – the first 5 shots should cover the space of answer labels.
Note that a few more than 5 shots per dataset were curated in the process, though in practice we are just using the first 5.
Appendix G OLMES prompt formats for each task
In Figure 5 we show an example of a full 5-shot prompt from ARC-Challenge (CF). Then we show single instance formatting for each of the 10 tasks in Figures 6- 25. For each task, we show both the MCF and CF formats.
All curated few-shot examples and prompt formatting code are available by accessing https://allenai.org/data/olmes.
Prompt Question: George wants to warm his hands quickly by rubbing them. Which skin surface will produce the most heat?
Answer: dry palms
Question: Which of the following statements best explains why magnets usually stick to a refrigerator door?
Answer: The refrigerator door contains iron.
Question: A fold observed in layers of sedimentary rock most likely resulted from the
Answer: converging of crustal plates.
Question: Which of these do scientists offer as the most recent explanation as to why many plants and animals died out at the end of the Mesozoic era?
Answer: impact of an asteroid created dust that blocked the sunlight
Question: Which of the following is a trait that a dog does NOT inherit from its parents?
Answer: the size of its appetite
Question: A boat is acted on by a river current flowing north and by wind blowing on its sails. The boat travels northeast. In which direction is the wind most likely applying force to the sails of the boat?
Answer:
Completion east
Figure 5: OLMES 5-shot prompt example for ARC-Challenge (CF).
Prompt Question: George wants to warm his hands quickly by rubbing them. Which skin surface will produce the most heat?
A. dry palms
B. wet palms
C. palms covered with oil
D. palms covered with lotion
Answer:
Completion A
Figure 6: OLMES prompt example for ARC-Challenge (MCF).
Prompt Question: George wants to warm his hands quickly by rubbing them. Which skin surface will produce the most heat?
Answer:
Completion dry palms
Figure 7: OLMES prompt example for ARC-Challenge (CF).
Prompt Question: Lichens are symbiotic organisms made of green algae and fungi. What do the green algae supply to the fungi in this symbiotic relationship?
A. carbon dioxide
B. food
C. protection
D. water
Answer:
Completion B
Figure 8: OLMES prompt example for ARC-Easy (MCF).
Prompt Question: Lichens are symbiotic organisms made of green algae and fungi. What do the green algae supply to the fungi in this symbiotic relationship?
Answer:
Completion food
Figure 9: OLMES prompt example for ARC-Easy (CF).
Prompt Persian language -- Persian, also known by its endonym Farsi, is one of the Western Iranian languages within the Indo-Iranian branch of the Indo-European language family. It is primarily spoken in Iran, Afghanistan (officially known as Dari since 1958), and Tajikistan (officially known as Tajiki since the Soviet era), and some other regions which historically were Persianate societies and considered part of Greater Iran. It is written in the Persian alphabet, a modified variant of the Arabic script, which itself evolved from the Aramaic alphabet.
Question: do iran and afghanistan speak the same language?
A. yes
B. no
Answer:
Completion A
Figure 10: OLMES prompt example for BoolQ (MCF).
Prompt Persian language -- Persian, also known by its endonym Farsi, is one of the Western Iranian languages within the Indo-Iranian branch of the Indo-European language family. It is primarily spoken in Iran, Afghanistan (officially known as Dari since 1958), and Tajikistan (officially known as Tajiki since the Soviet era), and some other regions which historically were Persianate societies and considered part of Greater Iran. It is written in the Persian alphabet, a modified variant of the Arabic script, which itself evolved from the Aramaic alphabet.
Question: do iran and afghanistan speak the same language?
Answer:
Completion yes
Figure 11: OLMES prompt example for BoolQ (CF).
Prompt Question: Sammy wanted to go to where the people were. Where might he go?
A. race track
B. populated areas
C. the desert
D. apartment
E. roadblock
Answer:
Completion B
Figure 12: OLMES prompt example for CommonsenseQA (MCF).
Prompt Question: Sammy wanted to go to where the people were. Where might he go?
Answer:
Completion populated areas
Figure 13: OLMES prompt example for CommonsenseQA (CF).
Prompt Health: How to cope with suicidal thoughts. Put off any plans. Promise yourself that you’ll wait 48 hours before doing anything. Remember, thoughts don’t have the power to force you to act.
Choose the best continuation:
A. Even when you do, there may be a small image of the future still lurking around your brain. For instance, don’t tell yourself that you can’t make it.
B. You’re doing something, and no one can force you to act. It’s completely natural to feel negative thoughts before you act.
C. Do not panic if people talk to you (even if it’s about quitting smoking). Have a plan for how you’re going to react to a group of people who bring on suicidal thoughts.
D. Sometimes extreme pain can distort our perception. Waiting before taking action will give your mind time to clear.
Answer:
Completion D
Figure 14: OLMES prompt example for HellaSwag (MCF).
Prompt Health: How to cope with suicidal thoughts. Put off any plans. Promise yourself that you’ll wait 48 hours before doing anything. Remember, thoughts don’t have the power to force you to act.
Completion Sometimes extreme pain can distort our perception. Waiting before taking action will give your mind time to clear.
Figure 15: OLMES prompt example for HellaSwag (CF).
Instruction The following are multiple choice questions (with answers) about abstract algebra.
Prompt Question: Find all c in Z_3 such that Z_3[x]/(x^2 + c) is a field.
A. 0
B. 1
C. 2
D. 3
Answer:
Completion B
Figure 16: OLMES prompt example for MMLU (abstract_algebra) (MCF).
Instruction The following are multiple choice questions (with answers) about abstract algebra.
Prompt Question: Find all c in Z_3 such that Z_3[x]/(x^2 + c) is a field.
Answer:
Completion 1
Figure 17: OLMES prompt example for MMLU (abstract_algebra) (CF).
Prompt Question: When standing miles away from Mount Rushmore
A. the mountains seem very close
B. the mountains are boring
C. the mountains look the same as from up close
D. the mountains seem smaller than in photographs
Answer:
Completion D
Figure 18: OLMES prompt example for OpenbookQA (MCF).
Prompt Question: When standing miles away from Mount Rushmore
Answer:
Completion the mountains seem smaller than in photographs
Figure 19: OLMES prompt example for OpenbookQA (CF).
Prompt Goal: how do you stab something?
A. stick a sharp object through it.
B. pin it with a sharp object.
Answer:
Completion A
Figure 20: OLMES prompt example for Physical Interaction QA (MCF).
Prompt Goal: how do you stab something?
Answer:
Completion stick a sharp object through it.
Figure 21: OLMES prompt example for Physical Interaction QA (CF).
Prompt Question: Cameron decided to have a barbecue and gathered her friends together. How would Others feel as a result?
A. like attending
B. like staying home
C. a good friend to have
Answer:
Completion A
Figure 22: OLMES prompt example for Social IQa (MCF).
Prompt Question: Cameron decided to have a barbecue and gathered her friends together. How would Others feel as a result?
Answer:
Completion like attending
Figure 23: OLMES prompt example for Social IQa (CF).
Prompt Fill in the blank: John moved the couch from the garage to the backyard to create space. The ___ is small.
A. garage
B. backyard
Answer:
Completion A
Figure 24: OLMES prompt example for WinoGrande (MCF).
Prompt1 John moved the couch from the garage to the backyard to create space. The garage
Prompt2 John moved the couch from the garage to the backyard to create space. The backyard
Completion is small.
Figure 25: OLMES prompt example for WinoGrande (CF). In this case the completions are the same for each answer choice, but the prompt is different.